# Unified Performance Control for Non-Square Nonlinear Systems with Relaxed Controllability

### v1

@@ III-B Error Transformation
It is seen from (10) that $s_{j}$ has the following properties: 1) for any $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, $s_{j}\rightarrow\pm\infty$ as $\zeta_{j}\rightarrow\bar{\delta}_{j}$ or $\zeta_{j}\rightarrow-\underline{\delta}_{j}$; 2) if $\zeta_{j}(0)$ satisfies $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$ and $s_{j}(t)$ is bounded $\forall t\geq 0$, then there exist some constants $-\underline{\delta}_{1j}$ and $\bar{\delta}_{1j}$ so that $-\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(0)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}$.
@@
[15] If the control to be developed is able to guarantee the boundedness of $s_{j}(t)$ for $t\geq 0$, then for any $\zeta_{j}(0)$ satisfying $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, (3) is ensured.
@@
The notable advantage of this error transformation is that the unified system framework for prescribed asymmetric and symmetric performance can be achieved by choosing the design parameters $\underline{\delta}_{j}$, $\bar{\delta}_{j}$ and $\varphi(0)$ (see [15] for details). Specifically, this error transformation framework can achieve global results similar to [22, 23, 24] (without considering error overshoot), as well as results similar to PPC [8, 20] and asymmetric performance behaviors.
III-C System Transformation
Denote ${\bm{\varepsilon}}_{1}=[{s}_{1},\cdots,{s}_{n}]^{T}$ as the transformed tracking error where ${s}_{j}$ ($j=1,\ldots,n$) is defined in (10). Differentiating (10) w.r.t. time, we have
@@ III-C System Transformation
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}({\bm{x}}_{1})+\bm{g}_{1}({\bm{x}}_{1},t)\bm{x}_{2}+{\bm{d}_{1}}({\bm{x}_{1}},t)-\dot{\bm{y}}_{d}+\bm{V}\right),$ (13)
where $\dot{\bm{\varepsilon}}_{1}=[\dot{s}_{1},\cdots,\dot{s}_{n}]^{T}$. According to the properties of the transformations (10), $\bm{W}$ and $\bm{V}$ are available for control design, moreover, $\bm{W}$ is invertible and $\bm{W}>0$ as long as $-\underline{\delta}_{j}<\zeta_{j}<\bar{\delta}_{j}$ over the interval $t\in[0,\tau_{\textrm{max}}]$. Now the original system (1) can be transformed as
$\dot{\bm{\varepsilon}}_{1}$ $=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}\bm{x}_{2}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)$
$\dot{\bm{x}}_{i}$ $=\bm{f}_{i}+\bm{g}_{i}\bm{x}_{i+1}+{\bm{d}_{i}},i=2,\ldots,N-1,$
$\dot{\bm{x}}_{N}$ $=\bm{f}_{N}+\bm{g}_{N}(\bm{\rho}\bm{u}+\bm{\upsilon})+{\bm{d}_{N}}.$ (14)
Therefore, (14) is still a strict-feedback system since $\bm{W}>0$, and hence, we can achieve the prescribed performance control of the original system (1) by designing the controller to stabilize the transformed system (14).
III-D Controller Design
To stabilize system (14), here we develop a robust adaptive controller by applying the backstepping method in [25]. Let us begin the control design by defining the virtual control errors as
${\bm{\varepsilon}}_{i}=\bm{x}_{i}-{\bm{a}}_{i-1},\ i=2,\ldots,N.$ (15)
where ${\bm{a}}_{i-1}$ is a virtual controller to be designed at the $i-1$th step and the actual controller $\bm{u}$ is derived at the final step.
Step 1: From (14) and (15), it is derived that
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}{\bm{\varepsilon}}_{2}+\bm{g}_{1}{\bm{a}}_{1}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right),$ (16)
Then, consider the Lyapunov candidate function as
$V_{11}=\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\bm{P}_{1}{\bm{\varepsilon}}_{1},$ (17)
@@ III-D Controller Design
${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}({\bm{f}}_{1}+{\bm{d}_{1}})\leq$ $\ \|\bm{W}\bm{\varepsilon}_{1}\|^{2}a_{11}^{2}a_{f1}^{2}\phi_{11}^{2}\phi_{f1}^{2}+\frac{1}{4},$
@@
with $\theta_{1}=\max\{a_{11}^{2}a_{f1}^{2},a_{11}^{2},a_{12}\}$ being an unknown positive constant and $\Phi_{1}=\phi_{11}^{2}\phi_{f1}^{2}+\phi_{11}^{2}(\|\dot{\bm{y}}_{d}\|+\|\bm{V}\|)^{2}+\frac{1}{2}(\lambda_{\textrm{min}}(\bm{W}))^{-2}\phi_{12}$ denoting a known and computable scalar function. Then, we have
@@
$\bm{a}_{1}$ $=-\kappa_{1}\bm{W}{\bm{\varepsilon}}_{1}-{\hat{\theta}_{1}{\Phi_{1}}{\bm{W}{\bm{\varepsilon}}_{1}}},$ (21)
$\dot{\hat{\theta}}_{1}$ $={\sigma_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}}\Phi_{1}-\mu_{1}\hat{\theta}_{1},\ \hat{\theta}_{1}(0)\geq 0,$ (22)
where $\kappa_{1}>0$, $\mu_{1}>0$ and $\sigma_{1}>0$ are design parameters, $\hat{\theta}_{1}$ is the estimate of unknown constant $\theta_{1}$. Now we consider the complete Lyapunov function candidate as
$V_{1}=V_{11}+\frac{1}{2\lambda_{1}\sigma_{1}}{\tilde{\theta}_{1}}^{2}.$ (23)
where $\tilde{\theta}_{1}=\theta_{1}-\lambda_{1}\hat{\theta}_{1}$ denotes the parameter estimate error. Differentiating (23) and using (20)-(22) yields
@@
$\bm{P}_{1}\bm{g}_{1}=\,\frac{\bm{G}_{1}}{2}+\frac{\left({\bm{P}_{1}\bm{g}_{1}-\bm{g}_{1}^{T}\bm{P}_{1}}\right)}{2},$ (25)
@@
Note that in (LABEL:sgs12), the potential destruction of the controllability condition by the performance incidental matrix $\bm{W}$ is neatly avoided by extending the ${\bm{\varepsilon}}_{1}$-based feedback to ($\bm{W}{\bm{\varepsilon}}_{1}$)-based feedback. Subsequently, inserting (25) and (LABEL:sgs12) into (24) yields
@@
$\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+\Delta_{1}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2},$
where the fact that $\frac{\mu_{1}}{\sigma_{1}}\tilde{\theta}_{1}\hat{\theta}_{1}\leq\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}$ is used, and $\Delta_{1}=\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}+\frac{1}{2}$ is a positive constant. The term ${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}$ will be handled in next step.
Step 2: The time derivative of ${\bm{\varepsilon}}_{2}=\bm{x}_{2}-\bm{a}_{1}$ is
$\dot{\bm{\varepsilon}}_{2}=\bm{f}_{2}+\bm{g}_{2}{\bm{\varepsilon}}_{3}+\bm{g}_{2}{\bm{a}}_{2}+{\bm{d}_{2}}-\dot{\bm{a}}_{1},$ (28)
@@
Choose the Lyapunov candidate function as $V_{21}=V_{1}+\frac{1}{2}{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}{\bm{\varepsilon}}_{2}$, where $\bm{P}_{2}\in\mathbb{R}^{n\times n}$ is defined in Assumption 3. Taking the time derivative of $V_{21}$ along (28) yields
@@
$\bm{a}_{2}$ $=-\kappa_{2}{\bm{\varepsilon}}_{2}-{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}},$ (32)
$\dot{\hat{\theta}}_{2}$ $={\sigma_{2}\|{\bm{\varepsilon}}_{2}\|^{2}}\Phi_{2}-\mu_{2}\hat{\theta}_{2},\ \hat{\theta}_{2}(0)\geq 0,$ (33)
@@
${\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}(\kappa_{2}{\bm{\varepsilon}}_{2}+{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}})\geq(\kappa_{2}\lambda_{2}+\lambda_{2}{\hat{\theta}_{2}{\Phi_{2}}})\|{\bm{\varepsilon}}_{2}\|^{2},$ (35)
@@
$+\Delta_{2}+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}.$
@@
$V_{i}=V_{i-1}+\frac{1}{2}{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}{\bm{\varepsilon}}_{i}+\frac{1}{2\lambda_{i}\sigma_{i}}{\tilde{\theta}_{i}}^{2},\ i=3,\ldots,N-1,$ (37)
@@
$\ \bm{a}_{i}=-\kappa_{i}{\bm{\varepsilon}}_{i}-{\hat{\theta}_{i}{\Phi_{i}}{{\bm{\varepsilon}}_{i}}},$ (38)
$\ \dot{\hat{\theta}}_{i}={\sigma_{i}\|{\bm{\varepsilon}}_{i}\|^{2}}\Phi_{i}-\mu_{i}\hat{\theta}_{i},\ \hat{\theta}_{i}(0)\geq 0,$ (39)
with $\Phi_{i}=\phi_{i1}^{2}\phi_{fi}^{2}+\phi_{i1}^{2}\phi_{f(i-1)}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\bm{\omega}_{i-1}\|^{2}+\phi_{(i-1)1}^{2}\|{\bm{\varepsilon}}_{i-1}\|^{2}+\frac{1}{2}\phi_{i2}$,
@@
$-\sum_{k=1}^{i}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}+\Delta_{i}+{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}\bm{g}_{i}{\bm{\varepsilon}}_{i+1},$
@@
$V_{N1}=V_{N-1}+\frac{1}{2}{\bm{\varepsilon}}_{N}^{T}\bm{P}_{N}{\bm{\varepsilon}}_{N},$ (42)
where $\bm{P}_{N}\in\mathbb{R}^{n\times n}$ is defined in Assumption 3.
Taking the time derivative of $V_{N1}$ along (41) , and from Assumption 3, it holds that
@@
$+\Delta_{N-1}+{\bm{\varepsilon}}_{N}^{T}\bm{\mathcal{G}}_{N}{\bm{u}}+\bm{h}_{N},$
@@
$\bm{u}$ $=-\frac{\bm{A}^{T}}{\|\bm{A}\|}\left(\kappa_{N}{\bm{\varepsilon}}_{N}+{\hat{\theta}_{N}{\Phi_{N}}{{\bm{\varepsilon}}_{N}}}\right),$ (46)
$\dot{\hat{\theta}}_{N}$ $={\sigma_{N}\|{\bm{\varepsilon}}_{N}\|^{2}}\Phi_{N}-\mu_{N}\hat{\theta}_{N},\ \hat{\theta}_{N}(0)\geq 0,$ (47)
@@
As shown in (21), (32), (38) and (46), the auxiliary matrix $\bm{P}_{i}$ is not used directly in the controller, here we just use $\bm{P}_{i}$ to construct the Lyapunov function $V_{i}$. It is worth emphasizing that by embedding some “core function”-based estimators in the backstepping design procedures, the system stability obstacles posed by the unknown auxiliary matrix $\bm{P}_{i}$ and nonlinear uncertainties are elegantly handled. Specifically, the constructed controllability conditions are satisfied online by adaptive matching $\bm{P}_{i}$, where neither an extra feasibility condition on $\bm{P}_{i}$ nor tedious offline solving of the matrix $\bm{L}$ is required, thus making the proposed controllability conditions more general than those in [6, 9, 13, 10, 20, 16].

### v2

@@ III-B Error Transformation
From (10), the variable $s_{j}$ exhibits the following characteristics: 1) for any $\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, $s_{j}\rightarrow\pm\infty$ as $\zeta_{j}\rightarrow\bar{\delta}_{j}$ or $\zeta_{j}\rightarrow-\underline{\delta}_{j}$; 2) if $\zeta_{j}(0)$ satisfies $\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$ and $s_{j}(t)$ remains bounded for all $t\geq 0$, there exist tighter bounds $\underline{\delta}_{1j}$ and $\bar{\delta}_{1j}$ within the original limits such that $\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(0)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}$.
@@
[15] If the control to be developed is able to guarantee the boundedness of $s_{j}(t)$ for $t\geq 0$, then for any $\zeta_{j}(0)$ satisfying $\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, (3) is ensured.
@@
The notable advantage of this error transformation is that the unified system framework for prescribed asymmetric and symmetric performance can be achieved by choosing the design parameters $\underline{\delta}_{j}$, $\bar{\delta}_{j}$ and $\varphi(0)$ (see [15] for details). Specifically, this error transformation framework can achieve global results similar to [28, 29, 30, 31] (without considering error overshoot), as well as results similar to PPC [8, 25] and asymmetric performance behaviors.
III-C Controller Design
To stabilize system (1) with unified prescribed performance, here we develop a robust adaptive controller by applying the backstepping method in [32]. Let us begin the control design by defining the transformed tracking error and the virtual control errors as
${\bm{\varepsilon}}_{1}$ $=[{s}_{1},\cdots,{s}_{n}]^{T}$ (11)
${\bm{\varepsilon}}_{i}$ $=\bm{x}_{i}-{\bm{a}}_{i-1},\ i=2,\ldots,N$ (12)
where ${s}_{j}$ ($j=1,\ldots,n$) is defined in (10), ${\bm{a}}_{i-1}$ is a virtual controller to be designed at the i1i-1th step and the actual controller $\bm{u}$ is derived at the final step.
Step 1: Differentiating (10) w.r.t. time, we have
@@ III-C Controller Design
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}\bm{x}_{2}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)$ (15)
where $\dot{\bm{\varepsilon}}_{1}=[\dot{s}_{1},\cdots,\dot{s}_{n}]^{T}$. According to (10), $\bm{W}$ and $\bm{V}$ are available for control design, moreover, $\bm{W}$ is invertible and $\bm{W}>0$ as long as $\underline{\delta}_{j}<\zeta_{j}<\bar{\delta}_{j}$ over the interval $t\in[0,\tau_{\textrm{max}}]$.
From (15) and (12), it is derived that
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}{\bm{\varepsilon}}_{2}+\bm{g}_{1}{\bm{a}}_{1}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)$ (16)
Next, consider the following Lyapunov candidate function:
$V_{11}=\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\bm{P}_{1}{\bm{\varepsilon}}_{1}$ (17)
@@
${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}({\bm{f}}_{1}+{\bm{d}_{1}})\leq$ $\ \|\bm{W}\bm{\varepsilon}_{1}\|^{2}a_{11}^{2}a_{f1}^{2}\phi_{11}^{2}\phi_{f1}^{2}+\frac{1}{4}$
@@
with $\theta_{1}=\max\{a_{11}^{2}a_{f1}^{2},a_{11}^{2},a_{12}\}$ being an unknown positive constant and $\Phi_{1}=\phi_{11}^{2}\phi_{f1}^{2}+\phi_{11}^{2}(\|\dot{\bm{y}}_{d}\|+\|\bm{V}\|)^{2}+\frac{1}{2}(\lambda_{\textrm{min}}(\bm{W}))^{-2}\phi_{12}$ denoting a known and computable scalar function. Note that in $\bm{h}_{1}$, the potential destruction of the controllability condition by the performance incidental matrix $\bm{W}$ is neatly avoided by extending the ${\bm{\varepsilon}}_{1}$-based feedback to ($\bm{W}{\bm{\varepsilon}}_{1}$)-based feedback.
Substituting (19) into (18), we have
@@
$\bm{a}_{1}$ $=-\kappa_{1}\bm{W}{\bm{\varepsilon}}_{1}-{\hat{\theta}_{1}{\Phi_{1}}{\bm{W}{\bm{\varepsilon}}_{1}}}$ (21)
$\dot{\hat{\theta}}_{1}$ $={\sigma_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}}\Phi_{1}-\mu_{1}\hat{\theta}_{1},\ \hat{\theta}_{1}(0)\geq 0$ (22)
where $\kappa_{1}>0$, $\mu_{1}>0$ and $\sigma_{1}>0$ are design parameters, $\hat{\theta}_{1}$ is the estimate of unknown constant $\theta_{1}$. We now introduce the complete Lyapunov function candidate as follows:
$V_{1}=V_{11}+\frac{1}{2\lambda_{1}\sigma_{1}}{\tilde{\theta}_{1}}^{2}$ (23)
where $\tilde{\theta}_{1}=\theta_{1}-\lambda_{1}\hat{\theta}_{1}$ is the parameter estimate error. Differentiating (23) and using (20)-(22) yields
@@
$\bm{P}_{1}\bm{g}_{1}=\,\frac{\bm{G}_{1}}{2}+\frac{\left({\bm{P}_{1}\bm{g}_{1}-\bm{g}_{1}^{T}\bm{P}_{1}}\right)}{2}$ (25)
@@
Subsequently, inserting (25) and (LABEL:sgs12) into (24) yields
@@
$\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+\Delta_{1}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}$
where the fact that $\frac{\mu_{1}}{\sigma_{1}}\tilde{\theta}_{1}\hat{\theta}_{1}\leq\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}$ is used, and $\Delta_{1}=\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}+\frac{1}{2}$ is a positive constant. The term ${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}$ will be addressed in the next step.
Step 2: Differentiating ${\bm{\varepsilon}}_{2}=\bm{x}_{2}-\bm{a}_{1}$ with respect to time, we obtain:
$\dot{\bm{\varepsilon}}_{2}=\bm{f}_{2}+\bm{g}_{2}{\bm{\varepsilon}}_{3}+\bm{g}_{2}{\bm{a}}_{2}+{\bm{d}_{2}}-\dot{\bm{a}}_{1}$ (28)
@@
Choose the Lyapunov candidate function as $V_{21}=V_{1}+\frac{1}{2}{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}{\bm{\varepsilon}}_{2}$, where the matrix $\bm{P}_{2}\in\mathbb{R}^{n\times n}$ is specified in Assumption 3. Taking the time derivative of $V_{21}$ along (28) yields
@@
$\bm{a}_{2}$ $=-\kappa_{2}{\bm{\varepsilon}}_{2}-{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}}$ (32)
$\dot{\hat{\theta}}_{2}$ $={\sigma_{2}\|{\bm{\varepsilon}}_{2}\|^{2}}\Phi_{2}-\mu_{2}\hat{\theta}_{2},\ \hat{\theta}_{2}(0)\geq 0$ (33)
@@
${\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}(\kappa_{2}{\bm{\varepsilon}}_{2}+{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}})\geq(\kappa_{2}\lambda_{2}+\lambda_{2}{\hat{\theta}_{2}{\Phi_{2}}})\|{\bm{\varepsilon}}_{2}\|^{2}$ (35)
@@
$+\Delta_{2}+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}$
@@
$V_{i}=V_{i-1}+\frac{1}{2}{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}{\bm{\varepsilon}}_{i}+\frac{1}{2\lambda_{i}\sigma_{i}}{\tilde{\theta}_{i}}^{2},\ i=3,\ldots,N-1$ (37)
@@
$\ \bm{a}_{i}=-\kappa_{i}{\bm{\varepsilon}}_{i}-{\hat{\theta}_{i}{\Phi_{i}}{{\bm{\varepsilon}}_{i}}}$ (38)
$\ \dot{\hat{\theta}}_{i}={\sigma_{i}\|{\bm{\varepsilon}}_{i}\|^{2}}\Phi_{i}-\mu_{i}\hat{\theta}_{i},\ \hat{\theta}_{i}(0)\geq 0$ (39)
with $\Phi_{i}=\phi_{i1}^{2}\phi_{fi}^{2}+\phi_{i1}^{2}\phi_{f(i-1)}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\bm{\omega}_{i-1}\|^{2}\\ +\phi_{(i-1)1}^{2}\|{\bm{\varepsilon}}_{i-1}\|^{2}+\frac{1}{2}\phi_{i2}$,
@@
$-\sum_{k=1}^{i}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}+\Delta_{i}+{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}\bm{g}_{i}{\bm{\varepsilon}}_{i+1}$
@@
$V_{N1}=V_{N-1}+\frac{1}{2}{\bm{\varepsilon}}_{N}^{T}\bm{P}_{N}{\bm{\varepsilon}}_{N}$ (42)
where $\bm{P}_{N}\in\mathbb{R}^{n\times n}$ is specified in Assumption 3.
Differentiating $V_{N1}$ with respect to time along (41) and applying Assumption 3, we obtain:
@@
$+\Delta_{N-1}+{\bm{\varepsilon}}_{N}^{T}\bm{\mathcal{G}}_{N}{\bm{u}}+\bm{h}_{N}$
@@
$\bm{u}$ $=-\frac{\bm{A}^{T}}{\|\bm{A}\|}\left(\kappa_{N}{\bm{\varepsilon}}_{N}+{\hat{\theta}_{N}{\Phi_{N}}{{\bm{\varepsilon}}_{N}}}\right)$ (46)
$\dot{\hat{\theta}}_{N}$ $={\sigma_{N}\|{\bm{\varepsilon}}_{N}\|^{2}}\Phi_{N}-\mu_{N}\hat{\theta}_{N},\ \hat{\theta}_{N}(0)\geq 0$ (47)
@@
Although the auxiliary matrix $\bm{P}_{i}$ appears in the Lyapunov function $V_{i}$, it is not directly used in the control laws (21), (32), (38) and (46). Instead, an adaptive mechanism based on “core function”-driven estimators is embedded into the backstepping design to implicitly handle the uncertainty in $\bm{P}_{i}$ and system nonlinearities. As a result, the proposed controllability conditions are satisfied online without requiring any feasibility assumptions on $\bm{P}_{i}$ or offline computation of auxiliary matrices such as $\bm{L}$, making the method more general than those in [6, 9, 13, 10, 25, 17].